OpenStudy (alli6106):

(a) At Davidson's Bike Rentals, it costs $33 to rent a bike for 8 hours. How many dollars does it cost per hour of bike use? (b) A color printer prints 24 pages in 8 minutes. How many minutes does it take per page? If necessary, round your answers to the nearest hundredth.

OpenStudy (alli6106):

I NEED HELP!!

OpenStudy (tkhunny):

Start answering. The first answer is given in the question. \("cost\;per\;hour"\longrightarrow \dfrac{Cost}{Hours} = \dfrac{$33}{8\;hr} = \dfrac{33}{8}\cdot\dfrac{$}{hr}\) Finish that up and tackle the second part.

OpenStudy (alli6106):

ok

OpenStudy (tkhunny):

Let's see what you get.

OpenStudy (alli6106):

i got 4.125

OpenStudy (tkhunny):

33/8 = 4.125 You seem to have misplaced your units. The correct answer is $4.125 / hour. Okay, do the next one.

OpenStudy (alli6106):

3?

OpenStudy (tkhunny):

Are you guessing? Prove it. Use notation that is helpful.

OpenStudy (alli6106):

idk

OpenStudy (tkhunny):

Hogwash. How did you get "3"?

OpenStudy (alli6106):

because 24/8 = 3?